# Completions of Kleene’s second model

### v5

@@ 5. Embeddings into $\mathcal{K}_{2}^{\kappa}$
The proof from [10] that every countable pca $\mathcal{A}$ embeds into $\mathcal{K}_{2}$ uses the recursion theorem, which we do not have available for large cardinalities, so in the proof below we have to avoid this. But we can still use the idea that because $|\mathcal{A}|\leqslant\kappa$, the graph of the application operator in $\mathcal{A}$ can be coded by an element of $\kappa^{\kappa}$.
@@
Referring to the coding before Definition 4.1, we define $f(a)$ by:
@@
Note that with this definition, the graph $\{(a,b,c)\mid\mathcal{A}\models a\cdot b\!\downarrow=c\}$ is coded into $f(a)$ for every $a$. Now we should check that (10) holds. Suppose that $a\cdot b\!\downarrow=c$ in $\mathcal{A}$. We check that Definition 4.1 gives $(f(a)\cdot f(b))(n)=f(c)(n)$ for every $n\in\kappa$. Suppose that $u,v,w$ are minimal codes of $a,b,c$, respectively.
@@
For $f(a)(\langle n\mbox{}^{\smallfrown}v\rangle)$ to produce the desired value, we need for (7) that both $f(a)(\langle n\rangle)$ and $f(a)(\langle\emptyset\rangle)$ are 00. Again, by assumption, $\langle\emptyset\rangle\notin L$, so $f(a)(\langle\emptyset\rangle)=0$. For every $n\in L$ we have $\langle n\rangle\notin L$, so indeed also $f(a)(\langle n\rangle)=0$ by (iv). ∎
@@ 6. Concluding remarks on completions
Asperti and Ciabattoni [1] proved that if Barendregt’s axiom (5) holds in a pca then it has a weak completion. (See also the discussion about this in Bethke et al. [5, p501].) By Engeler [6] (see footnote 1 above), the condition about Barendregt’s axiom is not needed. Note that Barendregt’s axiom does not hold for the example from Bethke et al. [5] of a pca without strong completions. Indeed, on p486 loc. cit. we have $a$ and $b$ distinct such that $s(sk)a=s(sk)b$, which violates Barendregt’s axiom. But since the example constructed in [5] is countable, and every countable pca weakly embeds into $\mathcal{K}_{2}$, it also follows from the strong completability of $\mathcal{K}_{2}$ that it has a weak completion.
Theorem 5.1 shows that, under GCH, every pca weakly embeds into a $\mathcal{K}_{2}^{\kappa}$ for suitably large $\kappa$. Since each $\mathcal{K}_{2}^{\kappa}$ is strongly completable, this has as a corollary that every pca has a weak completion (still assuming GCH). The latter statement follows unconditionally from Engeler [6]. GCH is needed in the proof of Theorem 5.1 to guarantee the existence of large cardinals $\kappa$ with $\kappa^{<\kappa}=\kappa$. However, we can get the corollary about completions unconditionally by using a binary version of $\mathcal{K}_{2}^{\kappa}$ defined on $2^{\kappa}$ rather than $\kappa^{\kappa}$. Namely, for the binary version we need the condition $2^{<\kappa}=\kappa$, and ZFC unconditionally proves the existence of arbitrary large $\kappa$ with this property. [^5] Namely, $2^{<\kappa}=\kappa$ holds for any $\kappa$ that is a strong limit, i.e. such that $\lambda<\kappa$ implies $2^{\lambda}<\kappa$, and the strong limit cardinals form a proper class, cf. Jech [11, p50]. This gives an alternative proof of Engeler’s result that every pca is weakly completable.
@@ Acknowledgments
We thank Henk Barendregt for pointing out that it follows from Engeler [6] that every pca has a weak completion, and for sharing the supplementary note Engeler [7]. We also acknowledge the helpful remarks of a referee that corrected an error in the proof of Lemma 4.3.

### v6

@@ 5. Embeddings into $\mathcal{K}_{2}^{\kappa}$
The proof from [GT23] that every countable pca $\mathcal{A}$ embeds into $\mathcal{K}_{2}$ uses the recursion theorem, which we do not have available for large cardinalities, so in the proof below we have to avoid this. But we can still use the idea that because $|\mathcal{A}|\leqslant\kappa$, the graph of the application operator in $\mathcal{A}$ can be coded by an element of $\kappa^{\kappa}$.
@@
Referring to the coding before Definition 4, we define $f(a)$ by:
@@
Note that with this definition, the graph $\{(a,b,c)\mid\mathcal{A}\models a\cdot b\!\downarrow=c\}$ is coded into $f(a)$ for every $a$. Now we should check that (10) holds. Suppose that $a\cdot b\!\downarrow=c$ in $\mathcal{A}$. We check that Definition 4 gives $(f(a)\cdot f(b))(n)=f(c)(n)$ for every $n\in\kappa$. Suppose that $u,v,w$ are minimal codes of $a,b,c$, respectively.
@@
For $f(a)(\langle n\mbox{}^{\smallfrown}v\rangle)$ to produce the desired value, we need for (7) that both $f(a)(\langle n\rangle)$ and $f(a)(\langle\emptyset\rangle)$ are 00. Again, by assumption, $\langle\emptyset\rangle\notin L$, so $f(a)(\langle\emptyset\rangle)=0$. For every $n\in L$ we have $\langle n\rangle\notin L$, so indeed also $f(a)(\langle n\rangle)=0$ by (iv).
@@ 6. Concluding remarks on completions
Asperti and Ciabattoni [AC97] proved that if Barendregt’s axiom (5) holds in a pca then it has a weak completion. (See also the discussion about this in Bethke et al. [BKdV99, p501].) By Engeler [Eng81] (see footnote 1 above), the condition about Barendregt’s axiom is not needed. Note that Barendregt’s axiom does not hold for the example from Bethke et al. [BKdV99] of a pca without strong completions. Indeed, on p486 loc. cit. we have $a$ and $b$ distinct such that $s(sk)a=s(sk)b$, which violates Barendregt’s axiom. But since the example constructed in [BKdV99] is countable, and every countable pca weakly embeds into $\mathcal{K}_{2}$, it also follows from the strong completability of $\mathcal{K}_{2}$ that it has a weak completion.
Theorem 6 shows that, under GCH, every pca weakly embeds into a $\mathcal{K}_{2}^{\kappa}$ for suitably large $\kappa$. Since each $\mathcal{K}_{2}^{\kappa}$ is strongly completable, this has as a corollary that every pca has a weak completion (still assuming GCH). The latter statement follows unconditionally from Engeler [Eng81]. GCH is needed in the proof of Theorem 6 to guarantee the existence of large cardinals $\kappa$ with $\kappa^{<\kappa}=\kappa$. However, we can get the corollary about completions unconditionally by using a binary version of $\mathcal{K}_{2}^{\kappa}$ defined on $2^{\kappa}$ rather than $\kappa^{\kappa}$. Namely, for the binary version we need the condition $2^{<\kappa}=\kappa$, and ZFC unconditionally proves the existence of arbitrary large $\kappa$ with this property. [^5] Namely, $2^{<\kappa}=\kappa$ holds for any $\kappa$ that is a strong limit, i.e. such that $\lambda<\kappa$ implies $2^{\lambda}<\kappa$, and the strong limit cardinals form a proper class, cf. Jech [Jec78, p50]. This gives an alternative proof of Engeler’s result that every pca is weakly completable.
@@ Acknowledgments
We thank Henk Barendregt for pointing out that it follows from Engeler [Eng81] that every pca has a weak completion, and for sharing the supplementary note Engeler [Eng21]. We also acknowledge the helpful remarks of a referee that corrected an error in the proof of Lemma 4.